The age of Mrs. Cruz is 9 years more than three times her son's age. Find the age of each if fifteen years from now, she will be 6 years more than twice her son's age
Answer by rothauserc(4718)

Let Mrs. Cruz's son's age be x and Mrs. Cruz's age by y
:
1) y = 3x +9
:
substitute for y in equation 2 using equation 1
:
2) y +15 = 2(x+15) +6
:
3x+9 +15 = 2x +30 +6
:
3x +24 = 2x +36
:
x = 12 and y = 45
:
*****************************************
Mrs. Cruz's son is 12 and Mrs. Cruz is 45
*****************************************
: